# Common Corpus: The Largest Collection of Ethical Data for LLM Pre-Training

### v1

@@ 2.1 Composition
The majority of the data in Common Corpus is in the public domain (see Table 3). The license for each document is provided in the metadata, so the dataset can be easily filtered by license as needed for a particular use case.
Common Corpus is multilingual (see Figure 1 and Table 3 for language coverage). Unlike many pretraining mixes, which are coposed mostly of English and code data, our dataset has strong representation for several languages, with the top nine languages constitute at least 10B tokens each [^3] Language distribution was computed using the fastText language identification model.. While we highlighted some of the issues faced in making open datasets above, all of these issues are exacerbated for languages other than English. Even in relatively high-resource languages like French, these problems are compounded by the fact that there is much less data available, and most tools generalize poorly to languages other than English. Additionally, Kreutzer et al. (2022) showed that many multilingual datasets contain a lot of low-quality or entirely unusable data. Many of the datasets they analyzed contained less than 50% of usable text, with 15 sources containing no usable data at all.
3 Provenance
In this section, we present the details about collections that comprise the Common Corpus, accompanied by the detailed information about the data sources, token count breakdowns, and the main included languages in Appendix B.
3.1 Open Government
Open Government is a set of financial, legal, and administrative data in the public domain. In total, the dataset contains more than 406B tokens and comprises two main datasets: Finance Commons and Legal Commons. See Appendix B.1 for detailed data composition.
Finance Commons.   This is the largest collection of financial documents in the public domain, comprising more than 14 billion words (more than 23 billion tokens). The documents come from a wide time range, from the mid-20th century all the way to 2024. Like many of our other datasets, Finance Commons is also multilingual. Most of the documents are in English, French, and German, but there are also texts in languages such as Romanian, Bulgarian, and Latvian. Additionally, this is a multimodal dataset. It includes more than 1.36 million original PDF documents from AMF and the WTO. The documents constitute a wide coverage of in-house layouts and formats produced by industrial and economic sectors. This makes this dataset ideal for developing the next generation of open-data multimodal models. One application for this dataset is to develop vision-language models (VLMs) for advanced document segmentation and processing. These documents also contain vast amounts of structured data, which is also a promising area of research that Finance Commons can help drive forward.
3.2 Open Culture
Open Culture is an aggregation of vast cultural heritage datasets containing both monographs and periodicals for over 13 languages: French, English, German, Spanish, Portuguese, Italian, Dutch, Luxembourgish, Danish, Swedish, Serbian, Czech, and Greek. There are also small portions of data in a wide variety of other languages, including Arabic, Bengali, Latin, Persian, Russian, Sanskrit, and Urdu.
A large part of Open Culture is compiled from Collections As Data (CAD) — large dumps of texts, datasets, PDFs, and even raw XML output (METS/ALTO). CAD initiatives, thus, considerably simplify dataset aggregation and are a major contribution to the digital commons ecosystem. All other parts of Open Culture have been collected on a resource-by-resource basis using APIs and other standard retrieval methods whenever available. The largest extractions of this kind include Internet Archive (about 2 million monographs in multiple languages) and Delpher (50,000 Dutch monographs and periodicals filtered to match the Dutch copyright law for public domain). We managed to compile a large multilingual collection despite such challenges, as poor OCR quality, which we partly solved through the development of OCR correction tools (see Section 4), text segmentation issues, and sometimes irrecoverable deterioration of the original support. For the detailed dataset composition, refer to Appendix B.2.
All Open Culture documents are in the public domain, which means their copyright has expired after a given term and there are no limitations on their reuse. For certain content, or in cases where we could not rely on the guarantee of established cultural heritage institutions, we implemented our own internal rights verification process. This process follows specific criteria, including author life and data object creation time, and takes into account that we only collected cultural heritage content from institutions based in the US or the EU (see the complete criteria list in Appendix C).
Open Culture data is also rich from a cultural and stylistic standpoint and can be used to train multilingual language models with more diverse and creative writing styles. As LLMs are trained on extremely large corpora to maximize next-word prediction accuracy, LLM-generated text can often lack in personality and be boring or generic (Jones and Bergen, 2024). This feature of language models stands in contrast with one of their most common uses. In an analysis of WildChat (Zhao et al., 2024), a dataset of 1 million user interactions with ChatGPT, Longpre et al. (2024c) found that over 30% of user requests involved creative compositions such as fictional stories, role-play, or poetry generation. At the same time, creative writing is poorly represented among datasets used to train LLMs, which mainly comprise web text (Longpre et al., 2024c). Therefore, Open Culture contributes data that can be used to train models for creative writing without violating copyright law. In addition, as many of the Open Culture datasets are historical (coming from the 18th-19th centuries, or even earlier; see Figure 2(a)), this collection also enables the development of historical language models. The metadata includes document creation year, which enables researchers to develop language models with a cutoff of the training data creation date.
3.3 Open Science
The Open Science collection includes scientific papers and other documents (theses, book reviews, clinical trials, etc). Following on the development of a global open access movement, these documents have been made increasingly available in open archives (preprints) or directly through open science publishers and infrastructure. Scientific content has become a primary focus for training data curation, due to its impact on performance on reasoning and advanced world knowledge tasks. Yet, the lack of licensing information has until now partly hindered reuse. The Semantic Scholar Open Research Corpus from Allen AI includes 81.1 million articles in English under an Open Data Commons Attribution License, allowing for the free reuse of the aggregated metadata while still acknowledging the remaining copyright of individual authors (Lo et al., 2020). The Pile incorporated dumps made available by arXiv and PubMed Central, also exclusively in English (Biderman et al., 2022). Finally, the BigScience project assembled several curated multilingual scientific datasets like the French HAL as part of the training data for BLOOM (Workshop et al., 2023).
The Open Science collection was made possible largely due to the recent development of OpenAlex [^4] https://openalex.org/, the largest open catalogue of scientific documents. OpenAlex maintains an expansive API search engine tracking detailed metadata for each indexed item, including the licensing, as well as a link to the original resource, which is generally in PDF format. We filtered OpenAlex on the three following licenses: CC-By, Public Domain/CC0, and CC-By-SA. The largest share of resources is available under CC-By, which is currently the recommended license by the Open Access definition. Open Science also includes smaller subsets, such as a direct extraction of arXiv articles available in CC-By and some European-specific resources not currently well indexed on OpenAlex (the exact distribution of token counts can be found in Appendix B.3).
3.4 Open Code
The Open Code collection comprises code data under a vast variety of free licenses, which allows NLP practitioners to train models on public domain code for either coding applications or in order to improve certain model performance on natural language reasoning, world knowledge tasks, mathematics, and structured output tasks (Aryabumi et al., 2024, Petty et al., 2024, MA et al., 2024). The code data we use comes from the Stack v1 and v2 (Kocetkov et al., 2023, Lozhkov et al., 2024). The Stack v1 contains 6.4TB of data and covers 30 programming languages, while the Stack v2 is approximately ten times bigger at 67.5TB and covers over 600 programming languages. All the code data is made available with a direct link to the original resource on GitHub. In total, Open Code contains 283,227,402,898 tokens (see most common languages in Appendix B.4).
[…]
3.5 Open Web
Wikipedia and Wikisource.   Wikimedia projects have always been major sources for language model training due to their reliability, extensive coverage, and textbook-like style. Despite this centrality, there is still a range of unresolved challenges with the most common versions available for training. The raw source of Wikimedia projects is made available in a specific mediawiki syntax, including a lot of project-specific models, tags, and conventions. The parsing of models is especially not straightforward, as they can either format existing text or remove or include external content (transclusion). As part of Wikimedia Enterprise, the Wikimedia Foundation created entirely new dumps from the rendered HTML sources, which in effect ensure that they include all the text made available to readers.
Youtube Commons.   YouTube Commons is our collection of audio transcripts of 2,063,066 videos uploaded on YouTube under a standardized CC-By license [^5] https://huggingface.co/datasets/PleIAs/YouTube-Commons.
A major objective will be the integration of web archives filtered by permissive license. Since 2016, several projects have attempted to reidentify Creative Commons licenses from web archives at scale including C4C (multilingual) (Habernal et al., 2016) and more recently CCCC (from Allen AI - in English) and most recently Common Crawl Creative Commons Corpus (C5; the first multilingual C4 derivative) [^6] https://huggingface.co/datasets/BramVanroy/CommonCrawl-CreativeCommons All theses projects struggled with license identification. While license mentions are frequently normalized with a direct link or logo to Creative Commons, there is no guarantee they really concern the entire content: “a blog page contains many photos and each photo is licensed under different CC-license type or a blog home page with many articles and each article is licensed under different CC-license type.”(Habernal et al., 2016). We hope this limitation could be overcome by a combination of web domain curation and fine-grained curation and annotation by a language model.
3.6 Open Semantic
Thanks to a collaboration with Wikimedia Deutschland, the entire set of Wikidata has been adapted in natural language and added to Common Corpus. This is to date the only available textual collection of Wikidata covering the entire range of 300 languages. Data processing involved the translation of items and properties into formal language sequences as simple natural language sequences, without textual synthesis: “Q41309 | P:27 | Q171150” becoming “Franz Liszt country of citizenship Kingdom of Hungary”. Within each entry, we provide all the available translations as consecutive blocks separated by a newline, anticipating that this may contribute to language alignment.
4 Cleaning and Curation
4.1 Text Segmentation
For text segmentation, we developed Segmentext [^7] https://huggingface.co/PleIAs/Segmentext (see example in Appendix D.1). It was trained to be resilient to broken and unstructured texts with digitization artifacts and ill-recognized layout formats. Given the diversity of the training data, Segmentext should work correctly on diverse document formats in the main European languages.
4.2 OCR Error Detection
We developed two pipelines to determine the digitization quality of different datasets in order to determine the appropriate treatment before datasets can be used for pre-training:
OCRoscope [^8] https://github.com/Pleias/OCRoscope.    This is a tool based on the language identification tool cld2999https://pypi.org/project/pycld2/. As cld2 cannot attribute non-recognized strings to any language, OCRoscope works by measuring the number of non-recognized 7-grams. OCRoscope provides a standardized rate of OCR quality and rate of non-character content for at least 80 languages. An example of OCPoscope work is presented in Appendix D.2. Complete processing by OCRoscope yielded a rate of 41% non-recognized 7-grams, which results in an OCR quality rate of 59%. In comparison, the self-estimated rate of OCR valid words by the French National Library is significantly higher (85%) for the whole document.
OCRerrcr [^10] https://huggingface.co/PleIAs/OCRerrcr.   This is a DeBERTa-v2-style language model with 400M parameters. OCRerrcr achieves higher accuracy than OCRoscope, but is more computationally intensive.
4.3 OCR Correction
We developed OCRonos [^11] https://huggingface.co/PleIAs/OCRonos model based on Llama 3 8B (Grattafiori et al., 2024). OCRonos is versatile and supports the correction of OCR errors, wrong word cuts or merges, and overall broken text structures. The training data includes a highly diverse set of OCR-ed texts in multiple languages, mostly coming from uncorrected versions of Open Culture and Open Government. On highly deteriorated content, OCRonos can act as a synthetic rewriting tool rather than a strict correction tool. An example of OCRonos work is presented in Appendix D.3. OCRonos contributes to making challenging resources usable for LLM applications and, more broadly, search retrieval. It is especially fitting in situations where the original PDF sources are too damaged for correct OCR or are not possible to retrieve.
OCRonos is generally faithful to the original material, provides sensible restitution of deteriorated text, and will rarely rewrite correct words. A common issue with OCR correction has been language switching: due to the inherent noise in the input text, an LLM will transcribe in a different language or script. The issue has been especially observed in smaller generalist models like GPT-3.5 or Claude-Haiku. OCRonos largely mitigates this issue.
4.4 PII Removal
Personally Identifiable Information (PII), which is any information that can be used to distinguish or trace an individual’s identity, is protected under legislation such as GDPR. Consequently, the new reguations put restrictions on LLM training data. In an analysis of large open datasets, there is a staggering amount of personal data in widely used datasets, e.g., large quantities of phone numbers in RedPajama, email addresses in S2ORC and peS2o, and IP addresses in the Stack (Elazar et al., 2024).
To identify and replace PII, we use Microsoft’s Presidio [^12] https://microsoft.github.io/presidio/, an open-source state-of-the-art tool. With the base settings, Presidio identified on average 55-60% of texts that included phone numbers due to different possible number formats. By applying custom regular expression patterns that include most phone numbers, we increased this accuracy to 85%. Typical methods of handling PII include removing it, replacing it with tags, and partial anonymization. These transformations substantially alter the format of PII, which could undermine the model’s understanding of the text or interfere with its ability to process text with real PII. Instead, we replace PII with fictuous but realistic values.
4.5 Toxicity Detection
In addition to posing legal and regulatory issues, web data is a major source of harmful and biased content (Common Crawl was shown to contain sexual content, hate speech, and racial and gender biases (Luccioni and Viviano, 2021)) and often suffers from low-quality and machine-generated text (Dodge et al., 2021). Public Domain data, e.g., in Open Culture, do not pose the same legal challenges but introduce new ones. Many texts there are historical periodicals and monographs from at least 80 years ago. Cultural norms surrounding the discussion of certain ethnic groups, women, and themes such as violence have changed dramatically. Many of these texts, therefore, do not meet modern ethical standards. Training models on these texts would lead to the reproduction and circulation of harmful language.
[…]
5 Impact
As early as August 2024, the first version of Common Corpus was highlighted as a “massive” part of pre-training data commons in European languages, especially contributing to making French a high-resource language (Ali and Pyysalo, 2024). In February 2025, Common Corpus was acknowledged as one of the main deliverables of the Paris AI Summit, as the “largest open database to train Large Language Models” [^15] https://www.elysee.fr/admin/upload/default/0001/17/373d6bcd8ea7f84a701c050bb9cab ccf3ed95c2b.pdf.
Common Corpus has already been incorporated in a wide range of LLM projects. As of May 2025, seven European SLMs and LLMs have been trained on at least some part of Common Corpus, and more will come throughout the year 2025. Among these are the models in the Pleias 1.0 family, including three SLMs (350 million, 1.2 billion, and 3 billion parameters), which were exclusively trained on Common Corpus. As such, they are fully reproducible in Europe. They were trained on the Jean Zay supercomputer in France (using H100 GPUs, under a Grand Challenge compute grant), and, specifically for the 1.2B model on the TractoAI’s cloud platform [^16] https://tracto.ai/. They come with a new tokenizer with 65,536 unique tokens, also trained on a representative sample of the Common Corpus for increased training efficiency and better multilingual support. During training, we experimented with different data mixtures of Common Corpus and derived a “refined” version for late training of about 1 trillion tokens, with more intensive, curation-based filtering. Two models derived from Pleias 1.0, Pleias-RAG-350M and Pleias-RAG-1B, are currently leading evaluations in their size range for retrieval augmented evaluation. Part of the performance can be attributed to the solid coverage of the main European language in Common Corpus, as well as a very extensive proportion of PDF-extracted sources that are very close to content commonly processed in the RAG production setting.
In January 2025, Barcelona Supercomputing Center released the Salamandra series of models under three different sizes: 2, 7, and 40 billion parameters (Gonzalez-Agirre et al., 2025). A major objective was enhanced support for a wide range of European languages, including lower-resource languages. The French part of Open Culture, French-PD, was incorporated into the overall data mixture. It was the main French dataset and the second-largest curated source. A French 7 billion parameters SLM, Lucie, incorporated a wider range of French sources from Common Corpus, not only the cultural heritage part but also some of the administrative sources (Gouvert et al., 2025). In addition, NeKo, a generalist language correction model trained by Nvidia, used the OCR-corrected part of Common Corpus both for training and evaluation (Lin et al., 2024). Anthropic also used Common Corpus in their experiments for feature visualization (Ameisen et al., 2025).
Common Corpus is also being used to create new datasets. For instance, YouTube Commons was used to create two major multimodal datasets: FineVideo (43000 YouTube videos) [^17] https://huggingface.co/datasets/HuggingFaceFV/finevideo and Mosel (1 million audio hours, half from YouTube Commons) (Gaido et al., 2024), which was used to train the FAMA series of speech foundational models (Papi et al., 2025).
6 Conclusion
Common Corpus is a significant contribution to the open LLM research community. Since the release of the first version, the data from Common Corpus has been used in numerous LLM projects by several European labs and industry leaders like Anthropic. Through the release of Common Corpus and this paper with thorough documentation of data collection and curation, we show that LLM development is possible while strictly adhering to the regulatory norms. We hope that Common Corpus will grow as a critical infrastructure for open science LLM research and development and inspire future initiatives in the open.
Limitations
Common Corpus is far from collecting the whole range of available open data, which we call the open data paradox. Continuing the work, we are currently collecting more data in European languages as part of the LLMs4EU project [^18] https://www.openaire.eu/llms4eu. In addition to this, even though the language diversity of the data is large, the coverage for low-resource languages is still shallow. We are developing in this direction, planning for enlargement of the corpora in African languages (e.g., Swahili, Wolof, Bambara). Furthermore, the amount of collected data (2 trillion tokens) is suitable for pre-training of models of limited size, as in the Pleias model family (see Section 5). Larger models, however will require significantly larger amounts of data. Therefore, future collection of permissible data is highly encouraged by this work.
In addition, Common Corpus naturally does not contain data for instruction-tuning and any forms of specialized tasks. Therefore, it is not directly suitable for task-specific fine-tuning. However, due to the multilingual, temporal, and semantic diversity of data, Common Corpus opens the opportunities for the creation of ethical fine-tuning datasets, as we did for the RAG models (see Section 5).
In Section 4, we described the tools we used for the data curation, filtering, and editing. Even though we used these methods responsibly and mitigated many issues overlooked by the counterparts (e.g., with toxicity detection and PII removal), none of the curation methods could naturally facilitate a hundred-percent accuracy. However, some issues, like OCR errors, present considerable challenges to the models and might even account for better handling of typos in the future. We would also like to mention that each data object is accompanied by sufficient metadata, and, if desired, LLM practitioners are free to filter out collections that might contain potential issues (as described in Section 4).
Acknowledgments and Disclosure of Funding
Appendix A Development
Common Corpus was first conceptualized in May 2023 as a French corpus for the French public services. The initial aim was to create a dataset of 100-140B tokens and use that data to train a reproducible small language model, modeled on the Pythia suite (Biderman et al., 2023). The initial nucleus of Common Corpus was French-PD a large collection of 85B words from monographs and newspapers in the public domain [^19] Released as https://huggingface.co/datasets/PleIAs/French-PD-Books and https://huggingface.co/datasets/PleIAs/French-PD-Newspapers..
A collection of various open datasets was first published under the name Common Corpus in March 2024202020Release blogpost: https://huggingface.co/blog/Pclanglais/common-corpus. At the time, it was a multilingual expansion of French-PD, retaining the initial focus on cultural heritage data but containing now about 500B words in the public domain from a variety of open collections in English, German, French, Italian, Spanish and other European languages. Throughout 2024, additional official releases gradually expanded this initial pretraining commons beyond the public domain, including Youtube Commons (in April 2024) and Finance Commons (in July 2024).
In November 2024, Common Corpus became an unified corpus, through the convergence of newly collected datasets by PleIAs and pre-existing ones (especially Wikimedia Projects). It contained two trillon tokens with the current division in five collections [^21] Release blogpost: https://huggingface.co/blog/Pclanglais/two-trillion-tokens-open. The current version of Common Corpus, the one we describe in this paper, was released in February 2025 for the Paris AI Summit. Common Corpus will remain an evolving resource reflecting the need for a continuous open infrastructure rather than a fixed dataset, with HuggingFace ensuring a proper archiving of past versions for reproducibility.
A primary driver of Common Corpus has been what we term the open data paradox: most of the content available under free licenses or uncopyrighted is not integrated into the existing pretraining datasets. Since GPT-3, the pretraining data infrastructure has been extensively reliant on web archiving and the constraints that stem from it. which include the lack of available tools for PDF hosting and processing, or for diverse language detection.
The open data paradox had the effect of making the open pretraining ecosystem viable in the first place. In particular, the gradual filtering of web archives according to more and more ethical rules leaves only a tiny residual amount under open licenses (less than 2% of the original set).
Appendix B Provenance
B.1 Open Govenment
@@ B.1.1 Finance Commons
The datasets that make up Finance Commons [^22] https://huggingface.co/collections/PleIAs/finance-commons-66925e1095c7fa6e6828e26c are presented in Table 4. Here, we also present the provenance details for each of the parts of Finance Commons:
@@
  • World Trade Organization (WTO). This dataset comprises documents from WTO’s official Documents Online platform [^25] https://docs.wto.org/dol2fe/Pages/FE_Search/FE_S_S005.aspx. The documents cover the years 1995 to 2024. Documents are available in three official languages: English, French, and Spanish. Some documents are available in other languages, e.g., Chinese, Korean, Arabic, German, and Portuguese. Also released separately as WTO-PDF [^26] https://huggingface.co/datasets/PleIAs/WTO-PDF.
  • French Authority for Financial Market (AMF). This is a dataset of documents from the French Authority for Financial Market, or the Autorité des marchés financiers [^27] https://www.amf-france.org/en/news-publications/publications/open-data (AMF), which is an independent public authority that regulates the French market. The documents are primarily in French. Also released separately as AMF-PDF [^28] https://huggingface.co/datasets/PleIAs/AMF-PDF.
@@
  • General Agreement on Tariffs and Trade (GATT) Library. This dataset comprises documents from GATT, which was an organization that promoted international commerce and the reduction of trade barriers among member states. Public documents were made available by the General Council of the WTO in 2006303030https://www.wto.org/english/docs_e/gattdocs_e.htm. The documents span from January 1, 1946, to September 6, 1996. Most of the documents are in English, but there are also documents in French, Spanish, and other languages.
B.1.2 Legal Commons

### v2

@@ Appendix B Corpus Composition
In Table 5, we present the top-50 languages in Common Corpus by token count. The token counts are presented in terms of our BPE tokenizer used to train the models described in Section 6, which was trained on a representative subsample of Common Corpus. To verify that our tokenizer serves as a strong baseline for token counts, we show its fertility in Appendix C.
Language Documents Words Tokens
English 154,175,907 634,794,970,595 968,757,721,747
French 035,245,624 162,061,620,874 275,358,437,630
German 011,385,377 056,674,819,173 112,127,458,251
Spanish 006,530,094 026,215,767,271 046,514,142,421
Latin 002,367,110 016,189,444,325 036,031,591,540
Italian 003,804,052 013,207,129,356 024,681,637,575
Polish 002,640,613 005,086,555,167 012,146,688,669
Greek 000844,122 003,796,018,483 011,376,498,056
Portuguese 001,756,922 005,234,373,473 010,262,747,943
Russian 002,762,818 003,222,919,854 009,439,453,633
Dutch 003,206,382 003,791,928,728 008,058,934,080
Danish 002,270,459 002,840,121,206 006,941,827,931
Slovak 000683,174 002,320,831,403 005,148,967,838
Czech 000946,534 001,966,829,784 004,798,558,092
Indonesian 001,023,361 001,660,129,567 004,381,878,823
Estonian 000685,180 001,613,093,565 004,379,534,617
Hungarian 000888,780 001,527,981,866 004,110,878,972
Swedish 003,250,289 001,782,642,556 004,014,927,806
Finnish 000931,201 001,356,653,003 003,943,036,413
Maltese 000480,491 001,421,372,608 003,646,102,921
Bulgarian 000576,997 001,444,748,621 003,422,182,324
Lithuanian 000539,906 001,192,586,834 003,097,400,907
Romanian 000725,766 001,398,178,029 002,909,452,579
Japanese 001,409,956 000204,698,439 002,738,872,745
Arabic 001,291,439 000827,392,227 002,682,255,180
Slovenian 000504,492 001,107,467,643 002,602,943,380
Latvian 000364,503 000966,005,785 002,579,350,119
Ukrainian 001,378,390 000786,829,868 002,561,253,212
Croatian 000443,669 000984,577,685 002,400,762,140
Chinese 001,426,017 000329,649,628 002,238,230,225
Haitian Creole 000292,718 000762,225,686 001,420,886,048
Turkish 000572,633 000380,551,359 001,206,732,266
Cebuano 006,123,694 000598,016,557 001,080,404,897
Norwegian Nynorsk 001,026,745 000405,911,677 001,036,914,970
Irish 000139,815 000371,175,864 000958,688,647
Castilian 000050,785 000457,490,798 000893,080,621
Serbian 000698,037 000292,399,267 000822,519,328
Hebrew 000343,406 000227,869,199 000763,029,488
Catalan 000803,532 000368,144,006 000736,220,789
Korean 000653,460 000143,729,419 000677,716,397
Persian 000983,632 000209,078,446 000668,287,975
Vietnamese 001,296,789 000267,190,147 000593,203,801
Norwegian 000628,798 000252,133,513 000561,694,623
Armenian 000306,959 000109,852,686 000539,900,475
Hindi 000210,984 000127,830,222 000463,832,707
Yiddish 000059,158 000122,100,684 000463,526,449
Welsh 000309,573 000191,827,552 000438,133,529
Occitan 000288,195 000195,380,666 000357,267,635
Georgian 000172,532 000071,678,084 000349,884,144
Basque 000439,582 000139,077,600 000348,891,265
Table 5: Top-50 languages in Common Corpus by token count. Each language is presented with its number of documents, words, and tokens in the corpus. The rows are ordered by the token count.
Appendix C Tokenizer Details
Figure 5: Comparing the fertility of PleIAs tokenizer (marked as “Ours 1B”) and other language models for six languages. The data source for all languages is the devtest set of Flores200 (Costa-jussà et al., 2022).
In Figure 5, we show how our tokenizer with a vocabulary of size 65,536, trained on a representative subsample of Common Corpus, compares to other language model tokenizers. Our tokenizer is outperformed only by Gemma 3, which has a tokenizer four times larger.
Appendix D Provenance
D.1 Open Govenment
@@ D.1.1 Finance Commons
The datasets that make up Finance Commons are presented in Table 6. Here, we also present the provenance details for each of the parts of Finance Commons:
@@
  • World Trade Organization (WTO). This dataset comprises documents from WTO’s official Documents Online platform. The documents cover the years 1995 to 2024. Documents are available in three official languages: English, French, and Spanish. Some documents are available in other languages, e.g., Chinese, Korean, Arabic, German, and Portuguese. Also released separately as WTO-PDF.
  • French Authority for Financial Market (AMF). This is a dataset of documents from the French Authority for Financial Market, or the Autorité des marchés financiers [^8] https://www.amf-france.org/en/news-publications/publications/open-data (AMF), which is an independent public authority that regulates the French market. The documents are primarily in French. Also released separately as AMF-PDF.
@@
  • General Agreement on Tariffs and Trade (GATT) Library. This dataset comprises documents from GATT, which was an organization that promoted international commerce and the reduction of trade barriers among member states. Public documents were made available by the General Council of the WTO in 2006101010https://www.wto.org/english/docs_e/gattdocs_e.htm. The documents span from January 1, 1946, to September 6, 1996. Most of the documents are in English, but there are also documents in French, Spanish, and other languages.
D.1.2 Legal Commons